A hollow steel tube is 30 ft. in length, 18 in. in diameter. Determine volume.
Volume = pi * height * (R^2 - r^2) R = larger radius r = smaller radius
V = 3.1416 * 30 * .75 * .75 * r smaller radius not given.
V = 53.0145 cu. ft.
Not sure of my calculations.

Found 2 solutions by addingup, Alan3354:
Answer by addingup(3677)

Volume:
9^2*3.1416*360 = 91609.06 cubic inches; 91609.06/1728 = 53.014 cubic feet
.
I haven't looked at your calculations but they must be right because you got to the correct answer.

Answer by Alan3354(69443)

A hollow steel tube is 30 ft. in length, 18 in. in diameter. Determine volume.
Volume = pi * height * (R^2 - r^2) R = larger radius r = smaller radius
V = 3.1416 * 30 * .75 * .75 * r smaller radius not given.
V = 53.0145 cu. ft.
Not sure of my calculations.
=================
Vol = pi*30*(0.75^2 - r^2) cubic feet
= 30pi*(0.5625 - r^2) cf --- r < 0.75
=============
You can't get a number without knowing r
------------------
If you ignore the wall thickness,
Vol = pi*30*0.75^2 cubic feet = the number you have.